40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 5 21:42:28 2009, Yoichi, Update, PSL, My thoughts on ISS INMONPD_Spectrum_1-10Hz.pdf
    Reply  Fri Feb 6 09:52:35 2009, Kakeru, Update, PSL, Current shunt transfar function TF_CS_gain.pngTF_CS_phase.png
       Reply  Fri Feb 6 23:23:48 2009, Kakeru, Yoichi, Update, PSL, ISS is fixed Current_ShuntTF_gain.pngCurrent_ShuntTF_phase.png
          Reply  Mon Feb 9 16:02:42 2009, Yoichi, Update, PSL, PSL relative intensity noise RIN-13dB.pngRIN-21dB.png
             Reply  Tue Feb 10 23:36:25 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
                Reply  Wed Feb 11 07:28:25 2009, Yoichi, Update, PSL, PA current and laser output 
                   Reply  Wed Feb 11 23:51:53 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
Message ID: 1276     Entry time: Thu Feb 5 21:42:28 2009     Reply to this: 1277
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: My thoughts on ISS 

Today, I worked with Kakeru on ISS.

The problem is sort of elusive. Some time, the laser power looks fine, but after a while you may see many sharp drops in the power. Some times, the power drops happen so often that they look almost like an oscillation.

We made several measurements today and Kakeru is now putting the data together. Meanwhile, I will put my speculations on the ISS problem here.

The other day, Kakeru took the transfer function of the ISS feedback filter (he is supposed to post it soon). The filter shape itself has a large phase margin ( more than 50deg ?) at the lower UGF (~3Hz) if we assume the response of the current shunt to be flat. However, when we took the whole open loop transfer function of the ISS loop, the phase margin was only 20deg. This leads to the amplification of the intensity noise around the UGF. The attached plot is the spectrum of the ISS monitor PD. You can see a broad peak around 2.7Hz. In time series, this amplified intensity noise looks like semi-oscillation around this frequency.

Since it is very unlikely that the PD has a large phase advance at low frequencies, the additional phase advance has to be in the current shunt. We measured the response of the current shunt (see Kakeru's coming post). It had a slight high-pass shape below 100Hz (a few dB/dec). This high-pass response produces additional phase advance in the loop.

There seems to be no element to produce such a high-pass response in the current shunt circuit ( http://www.ligo.caltech.edu/docs/D/D040542-A1.pdf )

This Jamie's document shows a similar high-pass response of the current ( http://www.ligo.caltech.edu/docs/G/G030476-00.pdf  page 7 )

Now the question is what causes this high-pass response. Here is my very fishy hypothesis :-)

The PA output depends not only on the pump diode current but also on the mode matching with the NPRO beam, which can be changed by the thermal lensing. If the thermal lensing is in such a condition that an increase in the temperature would reduce the mode matching, then the temperature increase associated with a pump current increase could cancel the power increase. This thermal effect would be bigger at lower frequencies. Therefore, the intensity modulation efficiency decreases at lower frequencies (high-pass behavior). If this model is true, this could explain the elusiveness of the problem, as the cancellation amount depends on the operation point of the PA. 

To test this hypothesis, we can change the pump current level to see if the current shunt response changes. However, the PA current slider on the MEDM screen does not work (Rob told me it's been like this for a while). Also the front panel of the MOPA power supply does not work (Steve told me it's been like this for a while). We tried to connect to the MOPA power supply from a PC through RS-232C port, which did not work neither. We will try to fix the MEDM slider tomorrow.

Attachment 1: INMONPD_Spectrum_1-10Hz.pdf  6 kB  Uploaded Thu Feb 5 20:25:31 2009  | Show | Show all
ELOG V3.1.3-